# Beyond Instruction Following: Evaluating Inferential Rule Following of Large Language Models

### v2

@@ 4.2 Inferential Rules Are Helpful for the Reasoning of LLMs
  • Golden Rule. This setting prompts the LLMs with the golden rule (i.e. a relevant rule that should be applied to the question) together with the original question.
  • Few Rule. This setting prompts the LLMs with the golden rule and two random irrelevant rules together with the original question.
  • All Rule. This setting is similar to Few Rule while the number of irrelevant rules increases to 30. This setting simulates a scenario where users prompt the LLMs with all possible inferential rules in the tasks instead of the relevant rules retrieved based on the query.
All these rule settings are tested in a zero-shot manner. As shown in Figure 4, in most cases, LLMs enjoy great performance improvements while being prompted with one golden inferential rule (No Rule $\rightarrow$ Golden Rule). Nevertheless, as the number of irrelevant rules increases, LLMs will find it hard to trigger and leverage the golden rule and thus have a performance drop (Golden Rule $\rightarrow$ Few Rule $\rightarrow$ All Rule).
Besides, we find that by following inferential rules, LLMs have better performance improvements on tasks that require complex reasoning, such as CLUTRR and CAIL2018. On the commonsense reasoning tasks, as the LLMs have parametric knowledge, the performance improvements brought by following inferential rules are relatively slim. Moreover, we find that all LLMs fail to follow the inferential rules in the task of TheoremQA, which illustrates the defect of current LLMs that can not follow complex mathematical or physical rules.
@@ 4.3 LLMs Prefer Natural Language Rules than Formal Language Rules
To evaluate whether LLMs can follow formal language rules, we transform the natural language rules of each benchmark into the form of First-Order Logic (FOL) by executing deterministic functions or prompting ChatGPT (Appendix B). Then we compare the reasoning performances of LLMs which are prompted by different forms of inferential rules in both zero-shot All Rule and Few Rule settings.
As shown in Figure 5, in most cases, LLMs conduct reasoning better with natural language rules than formal language rules. This aligns with our intuition that LLMs are mostly pre-trained with natural language and thus the inferential rules expressed with natural language are closer to the pre-trained distributions of LLMs than the inferential rules expressed with formal language. Nevertheless, in most cases, LLMs can follow the formal language rules. This reveals the possibility of learning formal language rules from a symbolic reasoning engine and then using LLMs for neural inference.
Figure 7: The inferential rule-following performance of LLMs when following factual and counterfactual rules.
Figure 8: The failure cases of LLMs inferential rule-following are classified into two categories: Triggering Error and Execution Error, which stand for the cases where LLMs fail to trigger the correct rule and LLMs fail to execute the correct rule, respectively.
4.4 Chain of Thought Is Inadequate for LLMs to Apply Rules
@@
4.5 LLMs Struggle to Follow Counterfactual Rules
Although we have verified the effectiveness of the inferential rules, it is still unclear whether LLMs completely follow the given inferential rules or merely use their parametric knowledge. Therefore, we designed the scenario of counterfactual rule-following.
@@ 4.5 LLMs Struggle to Follow Counterfactual Rules
As shown in Figure 7, in most cases of both Golden Rule and Few Rule settings, LLMs have significant performance drops when following counterfactual rules, compared with following factual rules. These results indicate that the performance improvements brought by following rules are partly attributed to the parametric knowledge of LLMs, besides following inferential rules.
4.6 Behavioral Analysis of LLMs Following Rules
To understand why LLMs fail to follow the given inferential rules in the reasoning process, we made a behavioral analysis of LLMs in the failure cases of LLMs inferential rule-following. Specifically, we adopt the few-shot Few Rule settings for LLMs to follow the rule-applying demonstrations to apply the given inferential rules to the current question. We ordered the LLMs first to choose an inferential rule to follow and then reason with it. By parsing the output of LLMs we can classify the failure cases of LLMs inferential rule-following into two categories: Triggering Error and Execution Error. Triggering Error indicates that the LLMs choose an irrelevant rule for the current case and therefore lead to an incorrect reasoning result. Execution Error indicates that although LLMs have chosen the correct rule for the current case, they fail to draw the correct conclusion of rule body. To faithfully describe the inferential rule-following behavior of LLMs instead of being affected by the parametric knowledge of LLMs, we run the analysis under the counterfactual settings of the selected benchmarks.
From the results shown in Figure 8, we can tell that when tackling different tasks, LLMs exhibit different behaviors in following rules. While rules have a heavy head for triggering (e.g. in CLUTRR and CAIL2018, the rule head will be a series of relational hops among characters), the LLMs are likely to make Triggering Errors. While the rule head is easy and commonsensical (e.g. in SALAD and ULogic), but the conclusion of the rule body is ambiguous or confused (the counterfactual scenario), the LLMs are likely to make Execution Errors.
To avoid Triggering Errors in the scenario of rule-enhanced reasoning with RAG paradigm (§2.1), the rule retriever plays a crucial role. The Triggering Errors can be eliminated if the rule retriever only retrieved the golden rules. However, existing works often employ simple sparse retrievers such as BM25 Yang et al. (2023); Sun et al. (2023); Zhu et al. (2023), which greatly compromises the inferential rule-following performance of LLMs.
@@ 4.6 Behavioral Analysis of LLMs Following Rules
4.7 Rule Following Capabilities of LLMs
To make a comprehensive evaluation of the inferential rule-following capability of the LLMs, we categorize the experimental results in the previous sections into 5 dimensions: Executing Rules, Triggering Rules, Following Formal Rules, Applying Rules, and Following Counterfactual Rules. The details of these dimensions are shown in Appendix C.
@@ 4.7 Rule Following Capabilities of LLMs
Generally, LLMs are not very good at inferential rule-following. This may be attributed to the lack of training in inferential rule-following in the current LLMs. As Instruction Fine-Tuning (IFT) has been a standard step in the pipeline of training LLMs and thus ensures their strong instruction-following capability, in the next section, we propose a fine-tuning method to effectively further improve the inferential rule-following capabilities of LLMs.
Rule Setting Fine-tuned CLUTRR SALAD CAIL2018 No Rule base 22.90 20.12 34.94 IFT 53.82 75.84 80.12 Few Rule base 40.36 88.91 63.25 IRFT 89.50 99.00 90.96 Golden Rule base 55.25 99.67 78.92 IRFT 100.0 99.90 87.95
Table 1: The performances of base and fine-tuned Llama-2-7b-chat on three datasets of RuleBench. IFT stands for instruction fine-tuning, i.e. fine-tuning the LLMs only with the questions and answers. IRFT stands for Inferential Rule-Following Tuning, i.e. train the LLMs to infer the answers with both questions and rules.
4.8 Inferential Rule-Following Tuning
To further improve the inferential rule-following capabilities of LLMs, we propose Inferential inferential rule-following Tuning (IRFT). Compared with IFT, IRFT involves inferential rules as a part of the prompt. The inferential rules can be only the golden rule or the golden rule with a few randomly sampled noise rules. This orders the LLMs to learn to infer the answer not only by the parametric knowledge but also by triggering and executing the golden rule. The tuning objective can be formalized as:
@@ 4.8 Inferential Rule-Following Tuning
Where the $q,r,a$ stands for the question, the golden rule, and the answer from the training set, respectively. $r_{i}\sim U(R)$ stands for randomly sampling $n$ rules from the entire rule sets as the noise rules. Based on the training data in RuleBench, we constructed training data in the settings of No Rule for IFT, and Few Rule ($n=2$) & Golden Rule ($n=0$) for IRFT.
As shown in Table 1, our proposed IRFT further significantly improves the performances of LLMs in the inferential rule-following scenarios and greatly outperforms IFT. This indicates that IRFT can effectively teach the LLMs the capabilities of inferential rule-following.
@@
5 Conclusion
In this paper, We introduce inferential rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following. We then construct and propose a new benchmark, RuleBench, for evaluating the inferential rule-following capabilities of LLMs. Based on RuleBench, we conduct a series of experiments to evaluate the inferential rule-following capabilities of 8 State-of-The-Art LLMs from different perspectives. We categorize the inferential rule-following capability in 5 dimensions and provide some insights into improvements for LLMs toward a better inferential rule-following intelligent agent. Finally, we propose the Inferential Rule-Following Tuning (IRFT), which further improves the inferential rule-following capabilities of LLMs.
Limitations
Although the evaluation results in this paper have illustrated the preference of LLMs in following rules, we have not yet proposed an effective method to help LLMs follow the rules better in a fixed given setting.
@@ Appendix A Instructions vs Rules
Nevertheless, we can not confine rules to instructions, or even identify instructions with rules Ribes-Inesta (2000). Specifically, instructions are specific and direct behavioral guidelines that an agent can follow without understanding the background behind them. Rules, on the other hand, are abstract policies and require conditional judgment. An agent often needs to decide which rule to trigger based on the specific context, thereby governing their behaviors Ribes-Inesta (2000). Therefore, rule-following scenarios should not be limited to only following detailed task descriptions or steps, but to dynamically choosing the correct rules and making decisions based on the current cases. Instructions tell LLMs what to reason, while rules tell LLMs how to reason.
@@ Appendix B Details of Constructing RuleBench
Here are the details of constructing each benchmark in RuleBench. The prompts used in this process are shown in Figure 9,10,11,12,13.

### v4

@@ 4.2 Inferential Rules Are Helpful for the Reasoning of LLMs
  • Golden Rule. This setting prompts the LLMs with the golden rule (i.e. an inferential rule that should be applied to the question) together with the original question.
  • Few Rule. This setting prompts the LLMs with the golden rule and two random noise rules together with the original question.
  • All Rule. This setting is similar to Few Rule while the number of noise rules increases to 30. This setting simulates a scenario where users prompt the LLMs with all possible inferential rules in the tasks instead of the relevant rules retrieved based on the query.
All these rule settings are tested in a zero-shot manner. As shown in Figure 4, in most cases, LLMs enjoy great performance improvements while being prompted with one golden inferential rule (No Rule $\rightarrow$ Golden Rule). Nevertheless, as the number of noise rules increases, LLMs will find it hard to trigger and leverage the golden rule and thus have a performance drop (Golden Rule $\rightarrow$ Few Rule $\rightarrow$ All Rule).
Besides, we find that by following inferential rules, LLMs have better performance improvements on tasks that require complex reasoning, such as CLUTRR and CAIL2018. On the commonsense reasoning tasks, as the LLMs have parametric knowledge, the performance improvements brought by following inferential rules are relatively slim. Moreover, we find that all LLMs fail to follow the inferential rules in TheoremQA, which illustrates the defect of current LLMs that can not follow complex mathematical or physical rules.
@@ 4.3 LLMs Prefer Natural Language Rules than Formal Language Rules
To evaluate whether LLMs can follow formal language rules, we transform the natural language rules of each benchmark into the form of First-Order Logic (FOL) by executing deterministic functions or prompting ChatGPT (Appendix B). Then we compare the reasoning performances of LLMs which are prompted by different forms of inferential rules in both All Rule and Few Rule settings.
As shown in Table 1, in most cases, LLMs conduct reasoning better with natural language rules than formal language rules (except for TheoremQA and Yi-1.5-6B-chat). This aligns with our intuition that LLMs are mostly pre-trained with natural language and thus the inferential rules expressed with natural language are closer to the pre-trained distributions of LLMs than the inferential rules expressed with formal language. This confirms the research motivation and value of Symbol-LLM Xu et al. (2023). Nevertheless, in most cases, LLMs still can follow the formal language rules. This reveals the possibility of learning formal language rules from a symbolic reasoning engine and then using LLMs for neural inference.
Model Factuality CAIL2018 SALAD CLUTRR ULogic
Golden Few Golden Few Golden Few Golden Few
gpt-4o Factual 98.19 96.99 99.74 97.93 81.18 65.46 88.71 96.55
C.F. 97.14 97.14  8.22 68.42 37.50  5.53 86.87 88.80
Llama-2-7b-chat Factual 62.65 70.63 70.68 41.54 77.58 37.79 91.81 87.71
C.F. 79.25 67.21 26.06 52.12 71.37 13.93 67.35 53.61
Yi-1.5-6B-Chat Factual 98.19 95.18 87.15 66.27 78.72 37.02 98.55 97.71
C.F. 87.43 80.57 71.24 56.68 70.13 19.27 73.61 71.08
gpt-3.5-turbo Factual 99.40 98.19 99.74 99.48 95.49 63.83 97.11 96.75
C.F. 97.14 91.43 100.0 100.0 98.50 43.86 81.69 75.78
gpt-4-turbo Factual 99.40 98.19 100.0 97.98 82.70 63.49 98.58 98.31
C.F. 99.43 100.0 12.93 77.55 78.72 36.16 86.51 86.99
Meta-Llama-3-8B Factual 89.16 86.14 96.09 93.35 62.21 40.17 98.92 98.80
C.F. 62.29 86.29 100.0 94.00 39.50  6.49 79.64 75.90
Mistral-7B-Instruct Factual 98.19 88.55 90.88 88.19 50.38 35.88 97.17 93.36
C.F. 88.00 65.14 62.10 87.48 26.15 13.26 33.61 24.10
Table 2: The performance of LLMs on RuleBench when following factual and counterfactual (C.F.) rules.
4.4 Chain of Thought Is Inadequate for LLMs to Apply Inferential Rules
@@
4.5 LLMs Struggle to Follow Counterfactual Inferential Rules
Although we have verified the effectiveness of the inferential rules, it is still unclear whether LLMs strictly follow the given inferential rules or merely the rules activate their parametric knowledge. Therefore, we designed the scenario of counterfactual rule-following.
@@ 4.5 LLMs Struggle to Follow Counterfactual Inferential Rules
As shown in Table 2, in most cases of both Golden Rule and Few Rule settings, LLMs have significant performance drops when following counterfactual rules, compared with following factual rules. These results indicate that the performance improvements brought by following rules are partly attributed to the parametric knowledge of LLMs, besides following inferential rules.
4.6 Behavioral Analysis of LLMs Following Inferential Rules
To understand why LLMs fail to follow the given inferential rules in the reasoning process, we made a behavioral analysis of LLMs in the failure cases of LLMs inferential rule-following. Specifically, we adopt the few-shot Few Rule settings for LLMs to follow the rule-applying demonstrations to apply the given inferential rules to the current question. We ordered the LLMs first to choose an inferential rule to follow and then reason with it. By parsing the output of LLMs we can classify the failure cases of LLMs inferential rule-following into two categories: Triggering Error and Execution Error. Triggering Error indicates that the LLMs choose a noise rule for the current case and therefore lead to an incorrect reasoning result. Execution Error indicates that although LLMs have chosen the correct rule for the current case, they fail to draw the correct conclusion of rule body. To faithfully describe the inferential rule-following behavior of LLMs instead of being affected by the parametric knowledge of LLMs, we run the analysis under the counterfactual settings of the selected benchmarks.
From the results shown in Figure 6, we can tell that when tackling different tasks, LLMs exhibit different behaviors in following rules. While rules have a heavy head for triggering (e.g. in CLUTRR and CAIL2018, the rule head will be a series of relational hops among characters), the LLMs are likely to make Triggering Errors. While the rule head is easy and commonsensical (e.g. in SALAD and ULogic), but the conclusion of the rule body is ambiguous or confused (the counterfactual scenario), the LLMs are likely to make Execution Errors. A case study of these two types of failures can be found in Appendix E.
To avoid Triggering Errors in the scenario of rule-enhanced reasoning with RAG paradigm (§2.1), the rule retriever plays a crucial role. The Triggering Errors can be eliminated if the rule retriever only retrieved the golden rules. However, existing works mostly employ simple sparse retrievers such as BM25 Yang et al. (2023); Zhu et al. (2023), which greatly compromises the inferential rule-following performance of LLMs.
@@ 4.6 Behavioral Analysis of LLMs Following Inferential Rules
Figure 6: The proportion of Triggering Error and Execution Error produced by LLMs on RuleBench.
4.7 Inferential Rule-Following Capabilities of LLMs
To make a comprehensive evaluation of the inferential rule-following capability of the LLMs, we categorize the experimental results in the previous sections into 5 dimensions: Executing Rules, Triggering Rules, Following Formal Rules, Applying Rules, and Following Counterfactual Rules. The details of these dimensions are shown in Appendix F.
@@ 4.7 Inferential Rule-Following Capabilities of LLMs
Generally, LLMs are not very good at inferential rule-following. This may be attributed to the lack of training in inferential rule-following in the current LLMs. In the next section, we propose a fine-tuning method to effectively further improve the inferential rule-following capabilities of LLMs.
Dataset CLUTRR SALAD CAIL2018 base 40.36 88.91 63.25 IRFT 42.94 91.80 67.44 Dataset TheoremQA ULogic DEER base 35.52 79.64 52.38 IRFT 43.03 87.83 61.90
Table 3: The performances of base and fine-tuned Llama-2-7b-chat on all datasets of RuleBench. The fine-tuning data is constructed using StringGame’s OOD synthetic data. Tested in zero-shot Few Rule setting.
5 Inferential Rule-Following Tuning
To further improve the inferential rule-following capabilities of LLMs, we propose Inferential inferential rule-following Tuning (IRFT). IRFT involves a golden rule with a few randomly sampled noise rules in the prompt and therefore orders the LLMs to learn to trigger and execute the golden rule. The tuning objective can be formalized as:
@@ 5 Inferential Rule-Following Tuning
Where the $q,r,a$ stands for the question, the golden rule, and the answer from the training set, respectively. $r_{i}\sim U(R)$ stands for randomly sampling $n$ rules from the entire rule sets as the noise rules.
To thoroughly separate the rule-following ability from domain knowledge, we propose to use purely synthetic data to construct the corpus, StringGame, for IRFT. The details of StringGame can be found in Appendix G.
As shown in Table 3, after training with IRFT on StringGame, the LLM enjoys a performance improvement on all tasks of RuleBench. We also evaluate both the base LLM and tuned LLM on an instruction-following benchmark, InfoBench, where their accuracies are 74.36% and 74.49%, respectively. Based on these results, we believe that inferential rule-following is an abstract and fundamental capability. Through IRFT, this capability can be abstracted and learned from purely synthetic symbolic tasks, allowing generalization to real-world rule-following tasks. Meanwhile, IRFT does not affect their general instruction-following ability. We also test using in-domain data for IRFT (Appendix H), and the performance improvements brought by IRFT are more significant.
6 Conclusion
In this paper, we introduce inferential rule-following as a vital capability of LLMs and distinguish it from the previous labors on instruction-following. We then construct and propose a new benchmark, RuleBench, for evaluating the inferential rule-following capabilities of LLMs. Based on RuleBench, we conduct a series of experiments to evaluate the inferential rule-following capabilities of 7 open-source and closed-source LLMs from different perspectives. We categorize the inferential rule-following capability in 5 dimensions and provide some insights into improvements for LLMs toward a better inferential rule-following intelligent agent. Finally, we propose the Inferential Rule-Following Tuning (IRFT), which further improves the inferential rule-following capabilities of LLMs.
Limitations
@@ Appendix A Instructions vs Rules
Nevertheless, we can not confine rules to instructions, or even identify rules with instructions Ribes-Inesta (2000). Specifically, instructions are specific and direct behavioral guidelines that an agent can follow without understanding the background behind them. Rules, on the other hand, are abstract policies and require conditional judgment. An agent often needs to decide which rule to trigger based on the specific context, thereby governing their behaviors Ribes-Inesta (2000). Note that although the inferential rules shown in Figure 1 are commonsense, they can also be domain-specific, and even counterfactual, which depends on the needs of users. Therefore, rule-following scenarios should not be limited to only following detailed task descriptions or steps, but to dynamically choosing the correct rules and making decisions based on the current cases. Following (Fagin et al., 1992), we call this type of rule inferential rule and named the scenario considered in this paper LLM inferential rule-following.
@@ Appendix B Details of Constructing RuleBench
Here are the details of constructing each benchmark in RuleBench. The prompts used in this process are shown in Figure 10,11,12,13,14.
@@
Figure 7: An example of Triggering Error on CLUTRR.
Figure 8: An example of Execution Error on CLUTRR.
Appendix C Test-time Prompts
The prompts used during test-time are shown in Figure 15,16,17,18,19,20. Please note that golden rules don’t always appear first; they are mixed with noise rules in random order during testing. The orders in these figures are for illustrative purposes.
Appendix D Evaluation of Inferential Rule-following on Long-tail Instances
To verify if following inferential rules is more beneficial when encountering long-tail instances, we selected instances with labels that are among the least frequent labels while ensuring that their amount is no less than 10% of the total dataset in SALAD (As their labels are meaningful) to serve as long-tail instances.
As shown in Table 4, closed-source LLMs (gpt-4o, gpt-4-turbo, gpt-3.5-turbo), compared with the performance on overall instances, although they have performance drops when following only instructions (No Rule) on long-tail instances, still achieve competitive performance while following inferential rules (All Rule, Few Rule, Golden Rule). This indicates that for closed-source LLMs, the improvement brought by following inferential rules on long-tail instances is greater compared to regular instances. However, for open-source LLMs, we found that they experience significant performance drops when following only instructions, and strangely, the effectiveness decreases with an increase in the number of noise rules, opposite to their behavior on the overall instances (Section 4.2). This indicates that these open-source LLMs have poor inferential rule-following capabilities on the distributions of long-tail samples.
Appendix E Case Study of Behavioral Analysis
Dataset Setting gpt-4o gpt-4-turbo gpt-3.5-turbo Llama-2-7b-chat Meta-Llama-3-8B Mistral-7B-Instruct Yi-1.5-6B-Chat
SALAD No Rule 47.25 54.85 53.70 20.12 39.55 53.51 50.94
All Rule 95.19 94.81 80.64 72.59 82.28 82.36 75.73
Few Rule 95.35 94.70 99.24 88.91 77.92 83.43 80.44
Golden Rule 98.59 99.36 99.67 99.67 95.11 92.28 81.78
SALAD long-tail No Rule 40.90 39.97 23.15  7.10  6.64  7.10  6.94
All Rule 95.83 92.28 77.47 75.46 75.15 75.00 74.07
Few Rule 96.91 97.68 99.23 55.71 56.26 56.94 53.70
Golden Rule 98.92 99.69 99.69 28.64 28.95 29.06 29.41
Table 4: The inferential rule-following performance of LLMs on all instances and only long-tail instances of SALAD.
To better illustrate what happens when LLMs fail to follow the inferential rules, we show two cases in which LLMs experience the Triggering Error (Figure 7) and Execution Error (Figure 8), respectively.
In the case of Triggering Error, LLMs appear to trigger an noise rule and consequently draw the wrong answer. In the case of Execution Error, although LLMs have triggered the golden rule, however, they fail to apply the rule to the question correctly, therefore also drawing the wrong answer.
Appendix F Details of The Dimensions
  • Triggering Rules. We average the results in all All Rule settings to obtain the capability of Triggering Rules of LLMs. This capability indicates how much the LLMs can resist the interruption of noise rules and find the golden rule.
Appendix G Details of The StringGame
The StringGame aims to use symbolic execution to construct a purely synthetic rule-following training dataset. It involves a series of five key steps that encompass rule generation, rule implementation, rule execution, prompt construction, and CoT generation. This task is about triggering and following the golden rule among the given rules to make the correct calculations based on the given string. An example is shown in Figure 9.
Example of StringGame Input: YTNRCMTZTVEAVQEHVJHW Rules: 1. If the length of the string is even, then output the number of unique consonants in the string. 2. If the number of "G"s is equal to 3, then output the index of the first "G" (index begins from 0). 3. If the string starts with a vowel, then output the total number of letters in the string divided by 2. 4. If the string starts and ends with the same letter, then output the index (1-based) of the first occurrence of the letter "R". Output: Based on the first rule, the answer is 12.
Figure 9: An example of StringGame.
G.1 Rule Generation
In this phase, a rule base is first initialized with 8 manually created seed rules. While the size of the rule base is smaller than 1000, we iteratively sample 5 rules from the rule base as few-shot demonstrations and ask the LLM to generate more similar but diverse rules. The response of the LLM will then be parsed and the new rules will be appended to the rule base.
G.2 Rule Implementation
Then, with two rule-function demonstrations, we use the LLM to implement each rule in the rule base to a snippet of the Python function. After that, another LLM, serving as the judge, will verify whether the generated code matches the provided rule. If the generated code passes the LLM validation and Python syntax check, it will then be saved as rule-function pairs.
G.3 Rule Execution
After all rules have been implemented as functions, we randomly generate strings with upper-case letters as the input. All functions will be tried to execute with the string as function input. If the function successfully returns a number, the corresponding rule will be noted as a “golden rule” for this string (meanwhile the function output will be saved as the answer to the question), otherwise (return None or get a run-time error), this rule will be noted as a “noise rule” for this string.
G.4 Prompt Construction
Having obtained the input strings together with their golden rules and noise rules, for each string, we sample one golden rule and randomly 0-3 noise rules. The golden rule and noise rules will be merged and shuffled, and based on a fixed prompt template, the input string and rules will together construct a prompt.
G.5 CoT Generation
Finally, we leverage STaR Zelikman et al. (2022) to sample success CoT trajectories from a closed-source LLM, i.e. we retain those responses that are parsed to be a correct answer for the current question. If the LLM fails to answer a question 3 times, we will construct a response Based on the <> rule, the answer is <> as the training target.
Rule Setting Dataset CLUTRR SALAD CAIL2018 Golden Rule base 55.25 99.67 78.92 IRFT 100.0 99.90 87.95 Few Rule base 40.36 88.91 63.25 IRFT 89.50 99.00 90.96
Table 5: The performances of base and fine-tuned Llama-2-7b-chat on three datasets of RuleBench. The fine-tuning data is constructed using RuleBench’s in-domain training data.
Appendix H In-domain IRFT on RuleBench
We also try to leverage the in-domain data for IRFT. Based on the training data in RuleBench (only the CLUTRR, SALAD, and CAIL2018 datasets contain training data), we constructed inferential-rule following samples under Few Rule ($n=2$) and Golden Rule ($n=0$) settings for IRFT.
As shown in Table 5, our proposed IRFT further significantly improves the performances of LLMs in the inferential rule-following scenarios and greatly outperforms the base model in both Few Rule and Golden Rule settings.